# LORE: Jointly Learning The Intrinsic Dimensionality and Relative Similarity Structure from Ordinal Data

### v1

@@ 1 Introduction
Figure 1: LORE jointly learns both the intrinsic dimensionality and relative similarities by balancing dimensionality with similarity constraints: Other methods require the embedding dimension to be chosen in advance, making them less data driven and often suboptimal.
@@ 5 Results
In this section, we present five pieces of empirical evidence in support of our method’s claims. First, we outline our experimental setup, including baseline methods and the generative process for producing ordinal embedding (OE) tasks. Next, we show how to select regularization levels for LORE. We then benchmark LORE against standard baselines across key metrics, followed by a comparison on proxy large language model (LLM) generated perceptual spaces. Finally, we assess performance on real, crowdsourced triplet data involving human judgments and see that LORE’s learned axes have semantic meaning. Collectively, these results demonstrate that LORE is uniquely effective at jointly learning high quality ordinal embeddings with intrinsic rank recovery, a property no other existing OE method can.
@@ 5.1 Setup
We benchmark LORE primarily against (1) SOE and (2) FORTE. These methods represent the best performing direct and Gram matrix OE approaches, respectively, as established by prior work (Vankadara et al., 2023). For our experiment on a simulated perceptual experiment and our last experiment on crowdsourced data, which is computationally less demanding, we also compare against additional OE methods like t-STE, CKL and the dimensionality estimation method using hypothesis testing from (Künstle et al., 2022) which is denoted as Dim-CV for simplicity. Our core evaluation criteria are:
@@
For LORE, we fix $p=0.5$ as it offers a good balance between rank recovery and optimization stability as seen in prior work (Lu et al., 2014; Wang et al., 2024; Sun et al., 2017). As $\mu$ only needs to be greater than the Lipschitz constant of triplet loss term, we set it to $0.1$ which is greater than what we find empirically. Both of these are set apriori and do not require tuning. The only hyperparameter that is tuned is the regularization parameter $\lambda$. However, as we show in Section 5.2, there exists a wide range of $\lambda$ around $0.01$ that yields both high triplet accuracy and intrinsic rank recovery across a variety of dataset conditions. Thus, we expect that in practice, a user can set $\lambda$ to $0.01$ without extensive hyperparameter tuning which our later experiments confirm works well across real world datasets. Initializations are usually critical for nonconvex optimization problems. However, due to the guaranteed convergence of LORE to stationary points, and the fact that Ordinal Embedding algorithms tend to have good local minima we find that random Gaussian initializations with variance of at least 5 work well across all experiments without any special tuning. All of our experiments use this initialization scheme and we suggest it as a default choice for practitioners.
@@ 5.4 LORE recovers Intrinsic Rank in a Simulated Perceptual Experiment
Human perceptual experiments are a key application of LORE, but for real perceptual datasets, the true intrinsic rank is unknown. Consequently, we leverage recent findings that large language models (LLMs) encode human-aligned perceptual information across domains such as taste, pitch, and timbre (Marjieh et al., 2024). Therefore we use the LLM embedding space as a realistic proxy of the true perceptual space in humans. To evaluate how our LORE compares to other baseline algorithms and the Dim-CV, which uses hypothesis testing and training multiple embeddings to estimate intrinsic rank, we evaluate both the test triplet accuracy and the measured rank of the learned embeddings. For Dim-CV we keep the range of hypothesed dimensions from 1 to 10 to match the true intrinsic ranks we test here.
In our experiment, we obtain SBERT embeddings (Reimers & Gurevych, 2019) for 50 randomly chosen foods, then control the intrinsic dimensionality for this experiment by applying truncated singular value decomposition (ranks 1-10). From this lower dimensional space, we generate noisy triplet comparisons (sampling 5% of the total, 30 repetitions per configuration, with noise 0.1) to mimic the data limited regime typical in human perceptual experiments. A detailed experimental setup is included in Appendix H.
@@ 5.5 LORE learns low rank and accurate representations on crowdsourced data
To test LORE in practical, noisy settings with unknown intrinsic rank, we evaluate it alongside baselines on three representative crowdsourced human similarity datasets covering food images (Wilber et al., 2014), material images (Lagunas et al., 2019), and car images (Kleindessner & Von Luxburg, 2017). These datasets differ in size, query semantics, and noise, reflecting the variety and challenges of real world ordinal data with results in Table 2. Further details are in Appendix E with additional real world dataset evaluations in Appendix J.
@@ 6 Discussion
Future directions include developing theoretical guarantees as well as exploring active learning methods to collect perceptual data more efficiently. Finally, we hope this work inspires further applied and theoretical advances, expanding the use of LORE for uncovering the structure of perceptual spaces across a range of domains.
@@
We have taken several steps to ensure the reproducibility of our work. All code and detailed documentation for our artificial human experiments and crowdsourced human experiments are included in the supplemental material, with full experimental configurations provided in Appendix H and Appendix I.
For the synthetic experiments and baseline comparisons, which require large scale parallelization and substantial computational resources, we do not provide raw code. Instead, we describe in detail the procedures and parameter settings necessary to reproduce them in Appendix G.
@@
We will release the full code and demo on GitHub and integrate the implementation of LORE into cblearn (Künstle & von Luxburg, 2024), a Python package for ordinal embeddings and comparison-based machine learning. This is forthcoming. We believe these efforts will make our work fully reproducible and accessible to the community.
@@ Acknowledgments
This work was supported by the National Science Foundation under Grant No. CCF-2107455. Sankar Alagapan was supported by the National Center for Advancing Translational Sciences of the National Institutes of Health under Award Numbers UL1TR002378 and KL2TR002381. Vivek Anand was supported by the National Institutes of Health under Data Science in Computational Neural Engineering Training Grant T32EB025816 and the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix F Additional Interpretability Plots
Appendix G Experimental Setup for Section 5.2 and Section 5.3
@@
  • Peak Signal to Noise Ratio: The PSNR is a measure of the quality of the recovered matrix. However, note that the recovered embedding matrix has to be aligned to the true percepts matrix to compute the PSNR. The specific formulation is described in Appendix M.
  • Normalized Procrustes Distance: The NPD is a measure of how well the recovered matrix matches the true matrix up to rotation, scaling and translation. To perform procrustes analysis, true percepts $\mathbf{P}\in\mathbb{R}^{N\times d}$ and the computed embedding $\mathbf{Z}\in\mathbb{R}^{N\times d^{\prime}}$ must be the same shape. Therefore, we use the same subspace alignment technique to ensure that the two matrices have the same shape. The specific formulation is described in Appendix M.
@@
Appendix H Experimental Setup for Section 5.4
@@
  • Peak Signal to Noise Ratio: The PSNR is a measure of the quality of the recovered matrix. However, note that the recovered embedding matrix has to be aligned to the true percepts matrix to compute the PSNR. The specific formulation is described in Appendix M. (We do not report these in the paper)
  • Normalized Procrustes Distance: The NPD is a measure of how well the recovered matrix matches the true matrix up to rotation, scaling and translation. To perform procrustes analysis, true percepts $\mathbf{P}\in\mathbb{R}^{N\times d}$ and the computed embedding $\mathbf{Z}\in\mathbb{R}^{N\times d^{\prime}}$ must be the same shape. Therefore, we use the same subspace alignment technique to ensure that the two matrices have the same shape. The specific formulation is described in Appendix M. (We do not report these in the paper)
@@
Appendix I Experimental Setup for Section 5.5
@@
Note that for this experiment unlike in Appendix H, we do not have access to the true percepts $\mathbf{P}$ and therefore cannot compute the PSNR or NPD. We only report the test triplet accuracy and measured rank. For Dim-CV we use a total of 5 cross validation folds but do not use multiple initializations due to computational constraints. As it is, Dim-CV is already two orders of magnitude slower than all of the OE algorithms due to the additional burden of training multiple embeddings due to the cross validation procedure. From our observations, increasing the number of cross validation folds and using multiple initializations reduces the standard deviation of the Dim-CV test triplet accuracy and rank but not the mean. Time however, increases linearly with the number of cross validation folds and initializations.
@@
Appendix J Additional Crowdsourced Dataset Experiments
@@
Appendix K Scalability of LORE
@@
Appendix L Convergence of LORE
@@
Taken together with Appendix K, these results indicate that LORE smoothly converges despite the non-convexity of the optimization problem and that the number of iterations taken to converge scales reasonably with dataset parameters. This empirically shows the provable convergence to a stationary point that we show in Appendix A. This result in combination with the fact that local minima for ordinal embedding problems are often good solutions (Vankadara et al., 2023) indicates that LORE is a scalable and practical algorithm for ordinal embedding in real world settings and that the theory supports the empirical findings.
Appendix M Formulation of Other Metrics
@@
M.2 Normalized Procrustes Distance
@@
M.3 Peak Signal to Noise Ratio

### v2

@@ 1 Introduction
Figure 1: LORE jointly learns both the intrinsic dimensionality and relative similarities by balancing dimensionality with similarity constraints: Other methods are susceptible to an incorrectly chosen embedding dimension.
@@ 5 Results
We present five pieces of empirical evidence in support of our method’s claims. We outline our experimental setup, including baseline methods and the generative process for producing ordinal embedding (OE) tasks. Next, we show how to select regularization levels for LORE. We then benchmark LORE against standard baselines across key metrics, followed by a comparison on proxy large language model (LLM) generated perceptual spaces. Finally, we assess performance on real, crowdsourced triplet data involving human judgments and see that LORE’s learned axes have semantic meaning. Collectively, these results demonstrate that LORE is the only OE method that can effectively jointly learn high quality ordinal embeddings and the intrinsic rank.
@@ 5.1 Setup
We benchmark LORE primarily against (1) SOE and (2) FORTE. These methods represent the best performing direct and Gram matrix OE approaches, respectively, as established by prior work (Vankadara et al., 2023). For our last two experiments, which are computationally less demanding, we also compare against additional OE methods like t-STE, CKL and the dimensionality estimation method using hypothesis testing from (Künstle et al., 2022) denoted as Dim-CV Our core evaluation criteria are:
@@
For LORE, we fix $p=0.5$ as it offers a good balance between rank recovery and optimization stability as seen in prior work (Lu et al., 2014; Wang et al., 2024; Sun et al., 2017). Our Results in Appendix G concur with the prior work. As $\mu$ only needs to be greater than the Lipschitz constant of triplet loss term, we set it to $0.1$ which is greater than what we find empirically. Both of these are set apriori and do not require tuning. The only hyperparameter that is tuned is the regularization parameter $\lambda$. However, as we show in Section 5.2, there exists a wide range of $\lambda$ around $0.01$ that yields both high triplet accuracy and intrinsic rank recovery across a variety of dataset conditions. Thus, we expect that in practice, a user can set $\lambda$ to $0.01$ without extensive hyperparameter tuning which our later experiments confirm works well across real world datasets. Initializations are usually critical for nonconvex optimization problems. However, due to the guaranteed convergence of LORE to stationary points, and the fact that Ordinal Embedding algorithms tend to have good local minima we find that random Gaussian initializations with variance of at least 5 work well across all experiments without any special tuning. All of our experiments use this initialization scheme and we suggest it as a default choice for practitioners.
@@ 5.4 LORE recovers Intrinsic Rank in a Simulated Perceptual Experiment
Human perceptual experiments are a key application of LORE, but for real perceptual datasets, the true intrinsic rank is unknown. Consequently, we leverage recent findings that large language models (LLMs) encode human-aligned perceptual information across domains such as taste, pitch, and timbre (Marjieh et al., 2024). Therefore we use the LLM embedding space as a realistic proxy of the true perceptual space in humans. To evaluate how our LORE compares to other baseline algorithms and the Dim-CV, which uses hypothesis testing and training multiple embeddings to estimate intrinsic rank, we evaluate both the test triplet accuracy and the measured rank of the learned embeddings. For Dim-CV we keep the range of hypothesed dimensions from 1 to 10 to match the true intrinsic ranks we test here. The measured rank is the averaged resulting rank of the embedding learned by each algorithm across 30 independent runs.
We first obtain SBERT embeddings (Reimers & Gurevych, 2019) for 50 randomly chosen foods, then control the intrinsic dimensionality for this experiment by applying truncated singular value decomposition (ranks 1-10). From this lower dimensional space, we generate noisy triplet comparisons (sampling 5% of the total, 30 runs, noise=0.1) to mimic the data limited regime typical in human perceptual experiments. A detailed experimental setup is included in Appendix I.
@@ 5.5 LORE learns low rank and accurate representations on crowdsourced data
To test LORE in practical, noisy settings with unknown intrinsic rank, we evaluate it alongside baselines on three representative crowdsourced human similarity datasets covering food images (Wilber et al., 2014), material images (Lagunas et al., 2019), and car images (Kleindessner & Von Luxburg, 2017). These datasets differ in size, query semantics, and noise, reflecting the variety and challenges of real world ordinal data with results in Table 2. Further details are in Appendix E with additional real world dataset evaluations in Appendix K.
@@ 6 Discussion
Future directions include developing theoretical guarantees and active learning methods to collect perceptual data more efficiently. We hope our work inspires further applied and theoretical advances applying LORE to uncover the structure of perceptual spaces across a range of domains.
@@
We have taken several steps to ensure the reproducibility of our work. All code and detailed documentation for our artificial human experiments and crowdsourced human experiments are included in the supplemental material, with full experimental configurations provided in Appendix I and Appendix J.
For the synthetic experiments and baseline comparisons, which require large scale parallelization and substantial computational resources, we do not provide raw code. Instead, we describe in detail the procedures and parameter settings necessary to reproduce them in Appendix H.
@@
The code to reproduce our results and a demo on how to use LORE are in https://github.com/vivek2000anand/lore_iclr. LORE’s integration into cblearn (Künstle & von Luxburg, 2024), a Python package for ordinal embeddings and comparison-based machine learning is forthcoming. We believe these efforts will make our work fully reproducible and accessible to the community.
@@ Acknowledgments
This work was supported by the National Science Foundation under Grant No. CCF-2107455. Additionally, supported in part by the NIH BRAIN Initiative through NIMH grant R61MH138966, the National Center for Advancing Translational Sciences of the National Institutes of Health under Award Number UL1TR002378 and KL2TR002381, and the Julian T. Hightower Chair at Georgia Tech. Vivek Anand was supported by the National Institutes of Health under the Data Science in Computational Neural Engineering Training Grant T32EB025816 and the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ Appendix F Additional Interpretability Plots
Appendix G Examining the effect of $p$ on LORE
In the following results we examine the effect of the $p$ in the ordinal embedding and rank recovery for LORE. Specifically, we vary $p$ with three values: 0.1, 0.5 (what we use in the paper) and 1 (equivalent to the nuclear norm). Like in C, we vary the number of percepts, noise and intrinsic rank while keeping the other parameters constant to see how $p$ affects the test triplet accuracy and rank recovery albeit at a reduced scale due to the computational cost of experiments. We include code for this experiment despite the additional time needed to run it. We find that $p=1$ has quite similar results to $p=0.5$ albeit with a reduced optimal regularization level of around $\lambda=10^{-2.5}$. However, we notice that the leeway to choose the regularization parameter is much higher in absolute terms for $p=0.5$ than $p=1$ as the performance of $p=1$ drops off more sharply as we deviate from the optimal regularization level. This is especially apparent for high noise levels. $p=0.1$ however is essentially useless to obtain lower rank solutions, at least for the range of regularization that we have examined. The baseline parameters we use for all of them, unless we are varying them, are number of percepts = 50, intrinsic rank = 5, fraction of queries = 0.1 and noise = 0.1.
Figure 16: Both $p=0.5$ and $p=1.0$ have stable rank recovery settings.. (Left) Mean test triplet accuracy vs $\lambda$ for LORE. (Right) Mean measured rank vs $\lambda$ for LORE.
Figure 17: $p=0.5$ has the most stable rank recovery settings as fraction of train triplets varies. (Left) Mean test triplet accuracy vs $\lambda$ for LORE as fraction of triplets used varies. (Right) Mean measured rank vs $\lambda$ for LORE as fraction of triplets used varies.
Figure 18: $p=0.5$ has the most stable rank recovery settings as number of percepts varies. (Left) Mean test triplet accuracy vs $\lambda$ for LORE as number of percepts varies. (Right) Mean measured rank vs $\lambda$ for LORE as number of percepts varies.
Figure 19: Both $p=0.5$ and $p=1.0$ have stable rank recovery settings as intrinsic rank varies. (Left) Mean test triplet accuracy vs $\lambda$ for LORE as intrinsic rank varies. (Right) Mean measured rank vs $\lambda$ for LORE as intrinsic rank varies.
Figure 20: Both $p=0.5$ and $p=1.0$ have stable rank recovery settings as noise varies. (Left) Mean test triplet accuracy vs $\lambda$ for LORE as noise varies. (Right) Mean measured rank vs $\lambda$ for LORE as noise varies.
Appendix H Experimental Setup for Section 5.2 and Section 5.3
@@ Appendix G Examining the effect of $p$ on LORE
  • Peak Signal to Noise Ratio: The PSNR is a measure of the quality of the recovered matrix. However, note that the recovered embedding matrix has to be aligned to the true percepts matrix to compute the PSNR. The specific formulation is described in Appendix N.
  • Normalized Procrustes Distance: The NPD is a measure of how well the recovered matrix matches the true matrix up to rotation, scaling and translation. To perform procrustes analysis, true percepts $\mathbf{P}\in\mathbb{R}^{N\times d}$ and the computed embedding $\mathbf{Z}\in\mathbb{R}^{N\times d^{\prime}}$ must be the same shape. Therefore, we use the same subspace alignment technique to ensure that the two matrices have the same shape. The specific formulation is described in Appendix N.
@@
Appendix I Experimental Setup for Section 5.4
@@
  • Peak Signal to Noise Ratio: The PSNR is a measure of the quality of the recovered matrix. However, note that the recovered embedding matrix has to be aligned to the true percepts matrix to compute the PSNR. The specific formulation is described in Appendix N. (We do not report these in the paper)
  • Normalized Procrustes Distance: The NPD is a measure of how well the recovered matrix matches the true matrix up to rotation, scaling and translation. To perform procrustes analysis, true percepts $\mathbf{P}\in\mathbb{R}^{N\times d}$ and the computed embedding $\mathbf{Z}\in\mathbb{R}^{N\times d^{\prime}}$ must be the same shape. Therefore, we use the same subspace alignment technique to ensure that the two matrices have the same shape. The specific formulation is described in Appendix N. (We do not report these in the paper)
@@
Appendix J Experimental Setup for Section 5.5
@@
Note that for this experiment unlike in Appendix I, we do not have access to the true percepts $\mathbf{P}$ and therefore cannot compute the PSNR or NPD. We only report the test triplet accuracy and measured rank. For Dim-CV we use a total of 5 cross validation folds but do not use multiple initializations due to computational constraints. As it is, Dim-CV is already two orders of magnitude slower than all of the OE algorithms due to the additional burden of training multiple embeddings due to the cross validation procedure. From our observations, increasing the number of cross validation folds and using multiple initializations reduces the standard deviation of the Dim-CV test triplet accuracy and rank but not the mean. Time however, increases linearly with the number of cross validation folds and initializations.
@@
Appendix K Additional Crowdsourced Dataset Experiments
@@
Appendix L Scalability of LORE
@@
Appendix M Convergence of LORE
@@
Taken together with Appendix L, these results indicate that LORE smoothly converges despite the non-convexity of the optimization problem and that the number of iterations taken to converge scales reasonably with dataset parameters. This empirically shows the provable convergence to a stationary point that we show in Appendix A. This result in combination with the fact that local minima for ordinal embedding problems are often good solutions (Vankadara et al., 2023) indicates that LORE is a scalable and practical algorithm for ordinal embedding in real world settings and that the theory supports the empirical findings.
Appendix N Formulation of Other Metrics
@@
N.2 Normalized Procrustes Distance
@@
N.3 Peak Signal to Noise Ratio